OpenStudy (anonymous):

What is the derivative of (4e)^x

OpenStudy (luigi0210):

Do you know what \(\large \frac{d}{dx}~a^x\) is?

OpenStudy (anonymous):

\[\huge y = (4e)^x \rightarrow y = 4^x e^x\] \[\huge \rightarrow y' = 4^x \frac{d}{dx} e^x + e^x \frac{d}{dx}4^x\] \[\huge \rightarrow y' = 4^x e^x + e^x 4^x \log 4\] \[\huge \rightarrow y' = 4^x e^x(1 + \log 4)\] is the required derivative. @abrandt